I oredered twice as many blankets than quilts. the total is 126. how many of each>
I am three years older than my sister. our combined age is 105. What is our ages?
bowling league has 16 teams in single ellimination. how many games does the winning team neeed to win?

Found 3 solutions by solver91311, rfer, unlockmath:
Answer by solver91311(24713)



Which part of "one question per post" did you not understand when the instructions for posting were presented?

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by rfer(16322)

126/3=42 quilts
84 blankets
--------------
x+3+x=105
2x=102
x=51
x+3=54

Answer by unlockmath(1688)

Hello,
I'll help you with the first one. Let's see if you can do the other two.
Let x represent the number of Quilts. We can now set up the following equation:
x+2x=126
Combine like terms and divide by 3 to get:
x=42
OK, There are 42 quilts and 84 blankets.
Make sense?
RJ
www.math-unlock.com